### v2

@@ Abstract
Multi-modal test-time adaptation (MM-TTA) is proposed to adapt models to an unlabeled target domain by leveraging the complementary multi-modal inputs in an online manner. Previous MM-TTA methods rely on predictions of cross-modal information in each input frame, while they ignore the fact that predictions of geometric neighborhoods within consecutive frames are highly correlated, leading to unstable predictions across time. To fulfill this gap, we propose ReLiable Spatial-temporal Voxels (Latte), an MM-TTA method that leverages reliable cross-modal spatial-temporal correspondences for multi-modal 3D segmentation. Motivated by the fact that reliable predictions should be consistent with their spatial-temporal correspondences, Latte aggregates consecutive frames in a slide window manner and constructs ST voxel to capture temporally local prediction consistency for each modality. After filtering out ST voxels with high ST entropy, Latte conducts cross-modal learning for each point and pixel by attending to those with reliable and consistent predictions among both spatial and temporal neighborhoods. Experimental results show that Latte achieves state-of-the-art performance on three different MM-TTA benchmarks compared to previous MM-TTA or TTA methods.
@@ 1 Introduction
Motivated by the above observation, we propose a novel MM-TTA method called ReLiable Spatial-temporal Voxels (Latte), which leverages the spatial-temporal correspondences within consecutive frames in a cross-modal manner as in Fig. 1b. Specifically, given frames of point clouds with their estimated poses, we regard points in the same voxel extracted from the merged point cloud frames as spatial-temporal correspondences. Different from previous works that merge all input frames [7, 5] or regularize frame-to-frame consistency [28], Latte aggregates consecutive frames in a sliding window manner to estimate the temporally local prediction consistency through the proposed Spatial-Temporal voxels (ST voxels) and entropy (ST entropy). Based on the reliability estimated by ST entropy in each modality, we conduct cross-modal learning in an adaptive attending manner to reduce the contributions of predictions from the noisy modality. To better facilitate the progress of MM-TTA, we combine our Latte with more recent backbones (SegFormer [43] and SPVCNN [33]) and conduct experiments on three different benchmarks, where Latte consistently outperform previous TTA or MM-TTA methods.
@@ 2 Related Works
Multi-modal domain adaptation for 3D segmentation. To avoid expensive annotation costs and overcome the poor generalizability of fully supervised solutions, various multi-modal domain adaptation methods have been investigated. Specifically for MM-UDA, xMUDA [16] is the primary works that explore incorporating cross-modal learning with MM-UDA. It regards cross-modal prediction consistency and pseudo-labels as its supervision signals in the unlabeled target domain. Most subsequent works propose solutions to address its limitations from different perspectives, such as more diverse point-pixel correspondence [24, 44], procedures to mitigate domain gaps [18, 20], and alleviating class-imbalanced problem [3]. Since MM-UDA is infeasible when facing online domain shifts, some previous works [29, 4] develop different methods to address the domain gap in an online manner, including cross-modal pseudo-label generation [29] and adaptive modality attending [4]. A recent work [30] proposes a source-free domain adaptation method by estimating cross-modal prediction agreement, yet it requires multiple training epochs and therefore is not suitable for online scenarios. Different from previous multi-modal domain adaptation methods that ignore temporal relationships, Latte couples spatial-temporal correspondence to mitigate the single-frame instability and achieve better online adaptation performance.
@@
Figure 2: Overall structure of Latte. Taking a student prediction frame of one modality as the query input, our slide-window extraction searches its spatial-temporal correspondences through voxelization within a time window so that temporally local prediction consistency can be established. ST voxels are then generated, where those with high ST entropy (larger than $\alpha$-quantile) are regarded as unreliable correspondences and discarded, while the others are leveraged for adaptive cross-modal learning by attending to the modality with lower ST entropy in a voxel-wise manner.
@@ 3.1 Frame-wise Predictions from Students and Teachers
Before exploiting spatial-temporal correspondence, we first generate the frame-wise predictions from each modality. Motivated by the fact that moving average models can provide more stable online predictions [34, 41, 4], Latte utilizes weight-averaged teacher models and fast student models for each modality as in previous works [29, 41], denoted as $\tilde{\phi}^{\textrm{m}}_{t}(\cdot)$ and $\phi_{t}^{\textrm{m}}(\cdot)$, respectively. Given each frame of 2D or 3D input $\textbf{{x}}^{\textrm{m}}_{t}$, the student predictions $\textbf{{p}}^{\textrm{m}}_{t}$ and teacher predictions $\tilde{\textbf{{p}}}^{\textrm{m}}_{t}$ are computed as:
@@ 3.2 Slide Window Frame Aggregation and Voxelization
where $\textbf{{T}}_{i}$ is the estimated pose at frame $i$ and so as $\textbf{{T}}_{j}$, which can be easily obtained online through off-the-shelf SLAM algorithms [36, 17]. $\textrm{cat}(\cdot)$ is the concatenation operation along the point number dimension while *** denotes the pose transformation. The voxelization is then performed on $\hat{\textbf{{x}}}_{i}^{\textrm{3D}}$, formulated as:
@@ 3.4 ST Voxel Aided Cross-Modal Learning
Additionally, the cross-modal attending is extended from the voxel level to the point level for better online predictions and cross-modal pseudo-label generation. Specifically, the ST entropy is propagated from the ST voxel to its reference point-wise prediction, indicating its confidence and consistency within its spatial-temporal neighborhoods. If one’s corresponding ST entropy has been filtered out as in Eq. (8), its ST entropy would fall back to the point-level entropy. Formally, for arbitrary point-wise prediction $\textbf{{p}}^{\textrm{m}}\in\textbf{{p}}_{\textrm{r}}^{\textrm{m}}$, the cross-modal predictions $y^{\textrm{xM}}$ are formulated as:
@@
     1. Get student and teacher predictions $\tilde{\textbf{{p}}}^{\textrm{m}}_{t},\textbf{{p}}^{\textrm{m}}_{t}$ by Eq. (1) for $i\in[t_{0}+1,t_{0}+B]$ do
         2. Get $\hat{\textbf{{x}}}_{i}^{\textrm{3D}},\textbf{{v}}^{\textrm{3D}}_{i},\textbf{{M}}^{\textrm{3D}}_{i}$ within the time window of size $w_{t}$ by Eq. (2)-(4) 3. Compute ST voxels $\textbf{{v}}_{i}^{\textrm{ST}}$ and ST entropy $E^{\textrm{m}}_{i}$ by Eq. (5)-(8) 4. Compute $\mathcal{L}^{\textrm{xM}}_{i}$ by Eq. (9)-(10)
@@
        5. Compute $\hat{E}_{\textrm{r}}^{\textrm{m}}$ and get $\textbf{{y}}^{\textrm{xM}}_{t}$ as predictions by Eq. (11)-(13) 6. Compute overall loss by Eq. (14) and update models by Eq. (15)
@@ 4.1 Benchmarks and settings
Implementation details. With the rapid progress in 2D and 3D segmentation, we employ all baseline methods and Latte with more recent SegFormer [43] (SegFormer-B1) and SPVCNN [33] (SPVCNN-cr1) as their 2D and 3D backbones, respectively, to investigate the improvement of existing TTA methods when combined with advanced networks. The pre-training procedures on the source domain follow a similar setting as in [16], except for 2D SegFormer which is trained with a base learning rate of 6E-5 with an AdamW optimizer and Poly scheduler as in [43]. For Latte, we maintain the same learning rate and optimizer as in the pre-training stage, while the scheduler is disabled. For voxelization, we leverage the off-the-shelf SLAM algorithm KISS-ICP [36] to generate poses and utilize a window size $w_{t}$ of 3 with a voxel size of 0.2, while the quantile $\alpha$ and coefficient $\lambda_{\textrm{xM}}$ are set to 0.9 and 0.3, respectively. $\lambda_{s}$ is empirically set to 0.99 as in [31, 4]. For all methods, we strictly follow the one-pass protocol to first evaluate the predictions of the input batch and then update the networks, where we only update the trainable parameters in normalization layers. For all baselines, we conduct a parameter search and report their best results for a fair comparison. We utilize the mean intersection over union (mIoU) as our evaluation metric. All experiments are conducted with PyTorch on a single RTX 3090. More details are presented in our appendix.
@@ 4.3 Ablation Studies and Visualziation
Table 2: Ablation studes about the effectiveness of cross-modal learning. When disabling $\textbf{{y}}^{\textrm{xM}}$, we utilize the average Softmax of all modalities as the pseudo-labels as a replacement. In experiment No. 8, we replace the ST entropy (ST ety) with point-wise entropy in Eq. (13) and set $w_{\textrm{v}}^{\textrm{m}}$ in Eq. (10) to 0.5. Here $\mathcal{L}^{\textrm{xM}}_{32}$ and $\mathcal{L}^{\textrm{xM}}_{23}$ refer to the former half and the latter half in Eq. 10, respectively.
@@
Figure 3: Ablation studies of different frame aggregation mechanisms. In addition to (a) Latte’s slide window aggregation, we test two different aggregation methods, including (b) replacing single frame student predictions with multiple frames in Eq. (5) and (c) adding non-overlapping aggregation to (b). Results show that our slide window aggregation can better evaluate the local consistency, leading to consistent performance improvement on all benchmarks.
@@
Figure 4: Parameter sensitivity analysis about various ST voxel size and confident percentile $\alpha$. All experiments are conducted across three benchmarks when other parameters are set to the default value.

### v3

@@ Abstract
Multi-modal test-time adaptation (MM-TTA) is proposed to adapt models to an unlabeled target domain by leveraging the complementary multi-modal inputs in an online manner. Previous MM-TTA methods for 3D segmentation rely on predictions of cross-modal information in each input frame, while they ignore the fact that predictions of geometric neighborhoods within consecutive frames are highly correlated, leading to unstable predictions across time. To fulfill this gap, we propose ReLiable Spatial-temporal Voxels (Latte), an MM-TTA method that leverages reliable cross-modal spatial-temporal correspondences for multi-modal 3D segmentation. Motivated by the fact that reliable predictions should be consistent with their spatial-temporal correspondences, Latte aggregates consecutive frames in a slide window manner and constructs Spatial-Temopral (ST) voxels to capture temporally local prediction consistency for each modality. After filtering out ST voxels with high ST entropy, Latte conducts cross-modal learning for each point and pixel by attending to those with reliable and consistent predictions among both spatial and temporal neighborhoods. Experimental results show that Latte achieves state-of-the-art performance on three different MM-TTA benchmarks compared to previous MM-TTA or TTA methods. Visit our project site https://sites.google.com/view/eccv24-latte.
@@ 1 Introduction
Motivated by the above observation, we propose a novel MM-TTA method called ReLiable Spatial-temporal Voxels (Latte), which leverages the spatial-temporal correspondences within consecutive frames in a cross-modal manner as in Fig. 1b. Specifically, given frames of point clouds with their estimated poses, we regard points in the same voxel extracted from the merged point cloud frames as spatial-temporal correspondences. Different from previous works that merge all input frames [7, 5] or regularize frame-to-frame consistency [28], Latte aggregates consecutive frames in a sliding window manner to estimate the temporally local prediction consistency through the proposed Spatial-Temporal voxels (ST voxels) and entropy (ST entropy). Based on the reliability estimated by ST entropy, we conduct cross-modal learning in an adaptive attending manner to reduce the contributions of predictions from the noisy modality. To better facilitate the progress of MM-TTA, we evaluate our Latte with more recent backbones (SegFormer [43] and SPVCNN [33]) on three different benchmarks, where Latte consistently outperforms previous TTA or MM-TTA methods.
@@ 2 Related Works
Multi-modal domain adaptation for 3D segmentation. To avoid expensive annotation costs and overcome the poor generalizability of fully supervised solutions, various multi-modal domain adaptation methods have been investigated. Specifically for MM-UDA, xMUDA [16] is the primary works that explore incorporating cross-modal learning with MM-UDA. It regards cross-modal prediction consistency and pseudo-labels as its supervision signals in the unlabeled target domain. Most subsequent works propose solutions to address its limitations from different perspectives, such as more diverse point-pixel correspondence [24, 44], procedures to mitigate domain gaps [18, 20], and alleviating class-imbalanced problem [3]. Since MM-UDA is infeasible when facing online domain shifts, some previous works [29, 4] develop different methods to address the domain gap in an online manner, including cross-modal pseudo-label generation [29] and adaptive modality attending [4]. A recent work [30] proposes a source-free domain adaptation method by estimating cross-modal prediction agreement, yet it requires multiple offline training epochs. Different from previous multi-modal domain adaptation methods that ignore temporal relationships, Latte couples spatial-temporal correspondence to mitigate the single-frame instability and achieve better online adaptation performance.
@@
Figure 2: Overall structure of Latte. Taking a student prediction frame of one modality as the query input, our slide-window extraction searches its spatial-temporal correspondences through voxelization within a time window to establish the temporally local prediction consistency. ST voxels are then generated, where those with high ST entropy (larger than $\alpha$-quantile) are discarded as unreliable correspondences, while the others are leveraged for adaptive cross-modal learning by attending to the modality with lower ST entropy in a voxel-wise manner.
@@ 3.1 Frame-wise Predictions from Students and Teachers
Before exploiting spatial-temporal correspondence, we first generate the frame-wise predictions from each modality. Motivated by the fact that moving average models can provide more stable online predictions [34, 41, 4], Latte utilizes weight-averaged teacher models and fast student models for each modality as in previous works [29, 41], denoted as $\tilde{\phi}^{\textrm{m}}_{t}(\cdot)$ and $\phi_{t}^{\textrm{m}}(\cdot)$, respectively. Given each input frame $\textbf{{x}}^{\textrm{m}}_{t}$, the student predictions $\textbf{{p}}^{\textrm{m}}_{t}$ and teacher predictions $\tilde{\textbf{{p}}}^{\textrm{m}}_{t}$ are computed as:
@@ 3.2 Slide Window Frame Aggregation and Voxelization
where $\textbf{{T}}_{i}$ is the estimated pose at frame $i$ and so as $\textbf{{T}}_{j}$, which can be easily obtained online through off-the-shelf SLAM algorithms [36, 17]. $\textrm{cat}(\cdot)$ is the concatenation operation along the point number dimension while * denotes the pose transformation. The voxelization is then performed on $\hat{\textbf{{x}}}_{i}^{\textrm{3D}}$, formulated as:
@@ 3.4 ST Voxel Aided Cross-Modal Learning
The cross-modal attending is further extended from the voxel level to the point level for better online predictions and cross-modal pseudo-label generation. Specifically, the ST entropy is propagated from the ST voxel to its reference point-wise prediction, indicating its confidence and consistency within its spatial-temporal neighborhoods. If one’s ST entropy has been filtered out as in Eq. (8), its ST entropy would fall back to the point-level entropy. For arbitrary point-wise predictions $\textbf{{p}}^{\textrm{m}}\in\textbf{{p}}_{\textrm{r}}^{\textrm{m}}$, the cross-modal predictions $y^{\textrm{xM}}$ are formulated as:
@@
     1. Get student and teacher predictions $\tilde{\textbf{{p}}}^{\textrm{m}}_{t},\textbf{{p}}^{\textrm{m}}_{t}$ by Eq. (1)
     for $i\in[t_{0}+1,t_{0}+B]$ do
         2. Get $\hat{\textbf{{x}}}_{i}^{\textrm{3D}},\textbf{{v}}^{\textrm{3D}}_{i},\textbf{{M}}^{\textrm{3D}}_{i}$ within the time window of size $w_{t}$ by Eq. (2)-(4)
         3. Compute ST voxels $\textbf{{v}}_{i}^{\textrm{ST}}$ and ST entropy $E^{\textrm{m}}_{i}$ by Eq. (5)-(8)
         4. Compute $\mathcal{L}^{\textrm{xM}}_{i}$ by Eq. (9)-(10)
@@
        5. Compute $\hat{E}_{\textrm{r}}^{\textrm{m}}$ and get $\textbf{{y}}^{\textrm{xM}}_{t}$ as predictions by Eq. (11)-(13)
         6. Compute overall loss by Eq. (14) and update models by Eq. (15)
@@ 4.1 Benchmarks and settings
Implementation details. With the rapid progress in 2D and 3D segmentation, we employ all baseline methods and Latte with more recent SegFormer [43] (SegFormer-B1) and SPVCNN [33] (SPVCNN-cr1) as their 2D and 3D backbones, respectively, to investigate the improvement of existing TTA methods when combined with advanced networks. The pre-training procedures on the source domain follow a similar setting as in [16], except for 2D SegFormer which is trained with a base learning rate of 6E-5 with an AdamW optimizer and Poly scheduler as in [43]. For Latte, we maintain the same learning rate and optimizer as in the pre-training stage, while the scheduler is disabled. For our voxelization, we leverage the off-the-shelf SLAM algorithm KISS-ICP [36] to generate poses and utilize a window size $w_{t}$ of 3 with a voxel size of 0.2m, while the quantile $\alpha$ and coefficient $\lambda_{\textrm{xM}}$ are set to 0.9 and 0.3, respectively. $\lambda_{s}$ is empirically set to 0.99 as in [31, 4]. For all methods, we strictly follow the one-pass protocol to first evaluate the predictions of the input batch and then update the networks, where we only update the trainable parameters in normalization layers. For all baselines, we conduct a parameter search and report their best results. We utilize the mean intersection over union (mIoU) as our evaluation metric. All experiments are conducted with PyTorch on a single RTX 3090. More details are presented in our appendix.
@@ 4.3 Ablation Studies and Visualziation
Table 2: Ablation studies about the effectiveness of cross-modal learning. When disabling $\textbf{{y}}^{\textrm{xM}}$, we utilize the average Softmax of all modalities as the pseudo-labels as a replacement. In experiment No. 8, we replace the ST entropy (ST ety) with point-wise entropy in Eq. (13) and set $w_{\textrm{v}}^{\textrm{m}}$ in Eq. (10) to 0.5. Here $\mathcal{L}^{\textrm{xM}}_{32}$ and $\mathcal{L}^{\textrm{xM}}_{23}$ refer to the former half and the latter half in Eq. 10, respectively.
@@
Figure 3: Ablation studies of different frame aggregation mechanisms. Besides (a) Latte’s slide window aggregation, we test two different aggregation methods, including (b) replacing single frame student predictions with multiple frames in Eq. (5) and (c) adding non-overlapping aggregation to (b). Results show that our slide window aggregation can better evaluate the local consistency and results in consistent improvement.
@@
Figure 4: Sensitivity analysis about the ST voxel size and confident percentile $\alpha$.
@@ 5 Conclusion
Acknowledgment
This research is supported by the National Research Foundation, Singapore, under the NRF Medium Sized Centre for Advanced Robotics Technology Innovation (CARTIN). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore.